OpenStudy (pagen13):

What is the slope-intercept for of the equation y-7=-5/2(x+4)

OpenStudy (shinalcantara):

slope-intercept form: \[y = mx + b\] where m= slope, b=y-intercept just rearrange the given equation such that the coefficient of y is one and in the form stated above

OpenStudy (pagen13):

so y=-5/2x+11?

jimthompson5910 (jim_thompson5910):

You probably made an error with distributing, but here is how I would do it \[\Large y-y_1=m(x-x_1)\] \[\Large y-7=-\frac{5}{2}(x+4)\] \[\Large y-7=-\frac{5}{2}x-\frac{5}{2}*4\] \[\Large y-7=-\frac{5}{2}x-10\] \[\Large y-7+7=-\frac{5}{2}x-10+7\] \[\Large y=-\frac{5}{2}x-3\]